# Is ChatGPT Transforming Academics’ Writing Style?

### v1

@@ 7 Discussion
The debate around the usage of generative models such as ChatGPT in academic writing is multi-faceted: from fears of lowering rigour due to “hallucinations” to uncertainty about the actual sources of AI-produced text. It is however indisputable that tools such as ChatGPT also have positive impacts: they help non-English native writers to improve the quality and flow of their text, as well as to translate into English from their mother tongue or vice versa. In this sense, generative AI is a great leveller, and as such it is a welcome addition to the academic’s toolbox. What we need to be wary of is its use in fully generative mode, without expert human supervision – something that we have not addressed in this paper.
We are aware that our methods can be further improved. For example, our results follow from analyzing a set of words selected based on the value of $q^{d}_{ij}$ and $\hat{r}_{ij}$. It is actually possible to fine-tune this criterium for a more accurate word selection, which would theoretically give better results, but would be more computationally expensive. Similarly, trying a larger range of prompts should theoretically result in better estimates. We are more interested in the density of ChatGPT style texts and its relative value (comparisons between categories and over time) than in establishing how many people are using ChatGPT – this can be estimated with the help of questionnaires, and it is not possible to get an accurate estimate only based on simulated data.
As our results have shown, ChatGPT is having an increasing impact on academic publications. This trend is hard to avoid, and we need to adapt gradually. With the increasing influx of young researchers, especially non-native English speakers, tools based on large language models, represented by ChatGPT, are transforming academic writing, at least for some disciplines. Even if you refuse to use them, you are likely to be influenced indirectly.
Appendix A Bonus!
Figure 11: Revise the following sentences:
Figure 12: Please rewrite the following paragraph from an academic paper:
Appendix B Period divisions
@@
Appendix C arXiv categories
@@
Appendix D Parameters
D.1 ChatGPT simulations
@@ D.2 Calibration
Appendix E Noise analysis
E.1 Variance in real data
Abstracts in the cs category among the first 500,000 articles were divided into groups in chronological order, with the same number in each group. We counted the number of occurrences of each word within each group, and calculated the variance between the different groups. This was repeated as a function of the number of abstracts included in each group, and the results are shown in Figure 13.
Figure 13: Variance of the word counts between groups of abstracts as a function of the number of abstracts included in each group.
Then we also analyzed the variance-to-mean ratio (defined as the variance of the sum of a word’s counts divided by the mean of the sum) and the coefficient of variation (defined as the standard deviation of the sum divided by the mean of the sum) for the 12 most frequent words, as shown in Figure 14 and Figure 15, and the variance-mean ratio of further words as in Figure 16.
Figure 14: Variance-to-mean ratio.
Figure 15: Coefficient of variation.
Figure 16: Variance-to-mean ratio for some further words.
@@ E.1 Variance in real data
E.2 Calculation details
@@
$\begin{split}&(\hat{\eta}^{g}_{j}(t)-\hat{\eta}_{j}(t))\sum_{i\in I_{j}}x^{2}_{ij}(t)\\ =&\sum_{i\in I_{j}}\frac{\partial g^{c}_{ij}(t)}{\partial\eta_{j}(t)}\left(h_{ij}(t)-\eta_{j}(t)x_{ij}(t)\right)\\ &-\sum_{i\in I_{j}}x_{ij}(t)g^{c}_{ij}(t)-\sum_{i\in I_{j}}g^{c}_{ij}(t)\frac{\partial g^{c}_{ij}(t)}{\partial\eta_{j}(t)}\end{split}$ (39)
@@
$\begin{split}\frac{\partial g^{c}_{ij}(t)}{\partial\eta_{j}(t)}=&f^{*}_{ij}(t)\epsilon^{\eta}_{ij}(q,f,t)+\eta_{j}(t)f^{*}_{ij}\frac{\partial\epsilon^{\eta}_{ij}(q,f,t)}{\partial\eta_{j}(t)}\\ &+\frac{\sqrt{f^{*}_{ij}(t)}}{2\sqrt{\eta_{j}(t)}}(\hat{r}_{ij}+\epsilon^{\eta}_{ij}(q,f,t))\delta_{ij}(1)\\ &+\sqrt{\eta_{j}(t)f^{*}_{ij}(t)}\frac{\partial\epsilon^{\eta}_{ij}(q,f,t)}{\partial\eta_{j}(t)}\delta_{ij}(1)\,.\end{split}$ (40)
@@
$\begin{split}\epsilon^{\eta}_{ij}(q,f,t)=&\frac{\epsilon_{ij}(1)}{\sqrt{\eta_{j}(t)f^{*}_{ij}(t)+\sqrt{\eta_{j}(t)f^{*}_{ij}(t)}\delta_{ij}(1)}}\\ &-\frac{\epsilon^{s}_{ij}(1)}{\sqrt{q^{d}_{ij}}}\end{split}$ (42)
@@
$\begin{split}&\frac{\partial\epsilon^{\eta}_{ij}(q,f,t)}{\partial\eta_{j}(t)}\\ =&\frac{-\left(2f^{*}_{ij}(t)\sqrt{\eta_{j}(t)}+\sqrt{f^{*}_{ij}(t)}\delta_{ij}(1)\right)\epsilon_{ij}(1)}{4\sqrt{\eta_{j}(t)}\left(\eta_{j}(t)f^{*}_{ij}(t)+\sqrt{\eta_{j}(t)f^{*}_{ij}(t)}\delta_{ij}(1)\right)^{\frac{3}{2}}}\,.\end{split}$ (43)
@@
Appendix F Other observations
We also define a change factor in the frequency of word $i$, ${R_{i}}^{\prime}$, as follows:
The total number of words in all abstracts of the first period is used as a base to normalize the frequency of words in the other periods, and the corresponding results are shown in Figure 17, Figure 18, and Figure 19.
Figure 17: The 12 words with the highest change rate ${R_{i}}^{\prime}$ and satisfying $\max_{t}({f_{i}}^{\prime}(t))>500$.
Figure 18: Examples of words with rapidly growing frequency in arXiv abstracts.
Figure 19: The words “are” and “is” are decreasing in frequency in arXiv abstracts.

### v2

@@ 7 Discussion
The debate around the usage of generative models such as ChatGPT in academic writing is multi-faceted: from fears of lowering rigour due to “hallucinations” to uncertainty about the actual sources of AI-produced text. It is however indisputable that LLM tools such as ChatGPT also have positive impacts: they help non-English native writers to improve the quality and flow of their text, as well as to translate into English from their mother tongue or vice versa. In this sense, generative AI is a great leveller, and as such it is a welcome addition to the academic’s toolbox. What we need to be wary of is its use in fully generative mode, without expert human supervision – something that we have not addressed in this paper.
We are aware that our methods can be further improved. For example, our results follow from analyzing a set of words selected based on the value of $q^{d}_{ij}$ and $\hat{r}_{ij}$. It is actually possible to fine-tune this criterium for a more accurate word selection, which would theoretically give better results, but would be more computationally expensive. Similarly, trying a larger range of prompts should theoretically result in better estimates. And better estimates may be made by more rigorous analysis, such as considering more complex noise terms. We are more interested in the density of LLM-style texts and its relative value (comparisons between categories and over time) than in establishing how many people are using LLMs – this can be estimated with the help of questionnaires, and it is not possible to get an accurate estimate only based on simulated data.
As our results have shown, that LLMs, started by ChatGPT, are having an increasing impact on academic publications. This trend is hard to avoid, and we need to adapt gradually. With the increasing influx of young researchers, especially non-native English speakers, LLM tools represented by ChatGPT, are transforming academic writing, at least for some disciplines. Even if you refuse to use them, you are likely to be influenced indirectly.
Appendix A Period divisions
@@
Appendix B arXiv categories
@@
Appendix C Other observations
We define the change factor in the frequency of word $i$, $R_{i}$, as follows:
$R_{i}=\frac{\max_{t}(f_{i}(t))-\min_{t}(f_{i}(t))}{\max_{t}(f_{i}(t))}$ (30)
where $f_{i}(t)$ is the count of word $i$ during the time period $t$.
Similarly, we define a change factor in the frequency of word $i$, ${R_{i}}^{\prime}$:
Figure 4(a) and Figure 4(b) illustrate that most of the words with the largest change rate in the time period considered (generally, an increase) in the abstracts are related to hot research topics of the last few years, such as “Covid-19”, “LLMs”, “AI”.
(a) The 12 words with the highest change rate $R_{i}$ and satisfying $\max_{t}(f_{i}(t))>500$.
(b) The 12 words with the highest change rate ${R_{i}}^{\prime}$ and satisfying $\max_{t}({f_{i}}^{\prime}(t))>500$.
Figure 4: Words with the highest change rate in frequency
The total number of words in all abstracts of the first period is used as a base to normalize the frequency of words in the other periods, and the corresponding results are shown Figure 5.
Figure 5: Word frequency changes (with different normalization) in abstracts.
Appendix D Correlation between simulated and real data
We also defined the word frequency change in all abstracts from year $t-1$ to year $t$, $R_{ij,t}$:
$R_{ij,t}=\frac{F_{ij,t}-F_{ij,t-1}}{F_{ij,t-1}}\,,$ (32)
Only words with a frequency larger than 0.1 times per abstract before ChatGPT processing are plotted in Figure 6(a) and Figure 6(b). The correlation coefficient between the word frequency change in arXiv abstracts and our estimated ChatGPT-induced word frequency change is very small in all four categories of abstracts, as shown in Figure 6(a).
(a) From 2021 to 2022.
(b) From 2022 to 2023.
Figure 6: Comparison of the predicted frequency change rate due to ChatGPT $\hat{r}_{ij}$ (x-axis) and the actual word frequency change for all abstracts (y-axis). CC indicates the correlation coefficient.
However, Figure 6(b) presents a totally different pattern, where $\hat{r}_{ij}$ and $R_{ij,2023}$ are strongly correlated, especially in computer science abstracts. Although many words seem insensitive to ChatGPT, we can still see a positive correlation for some words in this figure, even among the other categories.
Appendix E Parameters
E.1 ChatGPT simulations
@@ E.2 Calibration
Appendix F Noise analysis
F.1 Variance in real data
Abstracts in the cs category among the first 500,000 articles were divided into groups in chronological order, with the same number in each group. We counted the number of occurrences of each word within each group, and calculated the variance between the different groups. This was repeated as a function of the number of abstracts included in each group, and the results are shown in Figure 7(a).
(a) Variance of the word counts.
(b) Coefficient of variation.
Figure 7: Variance of the 12 most frequent words.
Then we also analyzed the coefficient of variation (defined as the standard deviation of the sum divided by the mean of the sum) for the 12 most frequent words, as shown in Figure 7(b), and the variance-to-mean ratio (defined as the variance of the sum of a word’s counts divided by the mean of the sum), as shown in Figure 8.
Figure 8: Variance-to-mean ratio
@@ F.1 Variance in real data
F.2 Calculation details
Case 1:
$g_{ij}(t)\frac{\partial g_{ij}(t)}{\partial\eta_{j}(t)}=g_{ij}(t)\frac{f^{*}_{ij}(t)(\eta_{j}(t)f^{*}_{ij}(t)+2\delta_{ij}(*))\epsilon_{ij}(1)}{2(\eta_{j}(t)f^{*}_{ij}(t)+\delta_{ij}(*))^{\frac{3}{2}}}-g_{ij}(t)\frac{f^{*}_{ij}\epsilon^{s}_{ij}(1)}{\sqrt{q^{d}_{ij}}}\,.$ (34)
Removing the items with zero means, we get
$\mathrm{E}\left[g_{ij}(t)\frac{\partial g_{ij}(t)}{\partial\eta_{j}(t)}\right]=\frac{\eta_{j}(t)(f^{*}_{ij}(t))^{2}(\eta_{j}(t)f^{*}_{ij}(t)+2\delta_{ij}(*))\sigma^{2}_{ij,\epsilon}}{2(\eta_{j}(t)f^{*}_{ij}(t)+\delta_{ij}(*))^{2}}+\frac{\eta_{j}(t)(f^{*}_{ij}(t))^{2}\sigma^{2}_{ij,\epsilon}}{q^{d}_{ij}}\,.$ (35)
Case 2:
We can define $g^{c}_{ij}(t)$ and $\xi^{c}_{ij}(t)$:
$\begin{split}g^{c}_{ij}(t)=&\eta_{j}(t)f^{*}_{ij}(t)\epsilon^{\eta}_{ij}(q,f,t)+\sqrt{\eta_{j}(t)f^{*}_{ij}(t)}(\hat{r}_{ij}+\epsilon^{\eta}_{ij}(q,f,t))\delta_{ij}(1)\end{split}$ (36)
$\xi^{c}_{ij}(t)=$ $\sqrt{f^{*}_{ij}(t)}\delta^{\prime}_{ij}(1)$ (37)
$L^{c}_{j,t,g}(\eta_{j})=\frac{1}{n_{j}}\sum_{i\in I_{j}}(h_{ij}(t)-\eta_{j}(t)x_{ij}(t)-g^{c}_{ij}(t))^{2}=\frac{1}{n_{j}}\sum_{i\in I_{j}}(\xi^{c}_{ij}(t))^{2}\,.$ (38)
And we will get a complex expression for the bias part like Eq. (23).
@@ F.2 Calculation details
$\begin{split}(\hat{\eta}^{g}_{j}(t)-\hat{\eta}_{j}(t))\sum_{i\in I_{j}}x^{2}_{ij}(t)=&\sum_{i\in I_{j}}\frac{\partial g^{c}_{ij}(t)}{\partial\eta_{j}(t)}\left(h_{ij}(t)-\eta_{j}(t)x_{ij}(t)\right)\\ &-\sum_{i\in I_{j}}x_{ij}(t)g^{c}_{ij}(t)-\sum_{i\in I_{j}}g^{c}_{ij}(t)\frac{\partial g^{c}_{ij}(t)}{\partial\eta_{j}(t)}\end{split}$ (39)
@@
$\begin{split}\frac{\partial g^{c}_{ij}(t)}{\partial\eta_{j}(t)}=&f^{*}_{ij}(t)\epsilon^{\eta}_{ij}(q,f,t)+\eta_{j}(t)f^{*}_{ij}\frac{\partial\epsilon^{\eta}_{ij}(q,f,t)}{\partial\eta_{j}(t)}+\frac{\sqrt{f^{*}_{ij}(t)}}{2\sqrt{\eta_{j}(t)}}(\hat{r}_{ij}+\epsilon^{\eta}_{ij}(q,f,t))\delta_{ij}(1)\\ &+\sqrt{\eta_{j}(t)f^{*}_{ij}(t)}\frac{\partial\epsilon^{\eta}_{ij}(q,f,t)}{\partial\eta_{j}(t)}\delta_{ij}(1)\,.\end{split}$ (40)
@@
$\epsilon^{\eta}_{ij}(q,f,t)=\frac{\epsilon_{ij}(1)}{\sqrt{\eta_{j}(t)f^{*}_{ij}(t)+\sqrt{\eta_{j}(t)f^{*}_{ij}(t)}\delta_{ij}(1)}}-\frac{\epsilon^{s}_{ij}(1)}{\sqrt{q^{d}_{ij}}}$ (42)
@@
$\frac{\partial\epsilon^{\eta}_{ij}(q,f,t)}{\partial\eta_{j}(t)}=\frac{-\left(2f^{*}_{ij}(t)\sqrt{\eta_{j}(t)}+\sqrt{f^{*}_{ij}(t)}\delta_{ij}(1)\right)\epsilon_{ij}(1)}{4\sqrt{\eta_{j}(t)}\left(\eta_{j}(t)f^{*}_{ij}(t)+\sqrt{\eta_{j}(t)f^{*}_{ij}(t)}\delta_{ij}(1)\right)^{\frac{3}{2}}}\,.$ (43)